OpenStudy (anonymous):

Write the rule of a function g whose graph can be obtained from the graph of the function f(x) = x2 - x + 3 by reflecting the graph in the x-axis, then shifting it vertically upward 5 units.

OpenStudy (anonymous):

g(x)= ???

OpenStudy (eyust707):

hehe im super rusty on these but can work you through it if oyu would like

OpenStudy (anonymous):

yes i would like lol

OpenStudy (eyust707):

ok so whats the grah look like?

OpenStudy (eyust707):

is it a line?

OpenStudy (eyust707):

a circle?

OpenStudy (anonymous):

no graph

OpenStudy (anonymous):

thats the only info they give me

OpenStudy (eyust707):

yes that is a function and has a graph

OpenStudy (anonymous):

its a parabola i think?

OpenStudy (eyust707):

yep!

OpenStudy (eyust707):

does it open up or down?

OpenStudy (anonymous):

we're doing transformations

OpenStudy (eyust707):

exactly we are trnsforming the graph

OpenStudy (anonymous):

up

OpenStudy (eyust707):



OpenStudy (eyust707):

what would we do to it to reflect it on the x axis?

OpenStudy (eyust707):



OpenStudy (eyust707):

we always want the function to be negative...

OpenStudy (anonymous):

so... shift to the right?

OpenStudy (eyust707):

well lets first reflect it then well pellet it

OpenStudy (anonymous):

ok proceed lol

OpenStudy (anonymous):

you there?

hero (hero):

I think it's g(x) = -x^2 + x + 2, but don't quote me

OpenStudy (eyust707):

my guess would be -x^2 + x + 8 -Hero heres my logic.. we want our y values to be 5 greater than what we got before so would we add 5 to the right side?

OpenStudy (anonymous):

Hero, you are correct. Tell me how you got your answer.

hero (hero):

Yes, I know I am correct

OpenStudy (eyust707):

ohh i seehehe gotta flip the whole thing

hero (hero):

Yup

OpenStudy (anonymous):

Please explain and draw out if you want as well

OpenStudy (eyust707):

hero want me to explain it?

hero (hero):

Go ahead

OpenStudy (eyust707):

ok so mario our original equation is x^2 - x + 3 = y we want to make all the +y valuse negitive so that it looks like the pic i drew above x^2 - x + 3 = -y that will flip it upside down

hero (hero):

What the...? C'mon man, you're confusing him

OpenStudy (eyust707):

multiply each side by -1 -x^2 +x -3 = y

OpenStudy (eyust707):

next we want to add 5 so that our y values are 5 greater

OpenStudy (eyust707):

-x^2 + x -3 +5 = y

OpenStudy (anonymous):

then...

OpenStudy (eyust707):

simplify.. thats my thought process anyways.. when i took this course i simply memorized what each transformation required.. it helps in the long run to tho to understand the graph and what must be done to it to change it

hero (hero):

There is something wrong with your steps and your approach

OpenStudy (eyust707):

whats that?

OpenStudy (anonymous):

Hero how did u do it?

OpenStudy (eyust707):

hehe it works in my mind lol us engineers think weird sometimes.. mario look in your book at the section that has transformations in it.. they describe it much better than i or hero

hero (hero):

I haven't even begun to describe it yet. I was allowing you to do it, but it didn't turn out so well.

OpenStudy (anonymous):

well begin then lol

OpenStudy (anonymous):

i got to go but when i come back tomorrow there better be an explanation lol

hero (hero):

Okay, so we have f(x) = x2 - x + 3 and we were asked to flip f(x) over the x axis, and then, shift it up vertically by 5. In other words, change f(x) to -f(x) + 5 Now, if we flip f(x) over the x axis, we will have -f(x), meaning the opposite of f(x). This simply means that all of the signs of f(x) will change. So -f(x) becomes -x^2 + x - 3 And then -f(x) + 5 becomes -x^2 + x + 2 Which is the correct answer.

OpenStudy (anonymous):

Ok that makes sense. Thanks.